three vectors from origin a,b,c.the ends of vec a is Q ,R,P respectively..if PQ=2QR then
(a)2a+c=3b (b) a+3c=2b (c) ba+b=2b (d) a+2c=3b

Grade:11th

1 Answers

Pushkar Aditya
10 years ago
c+PR=b (1) b+RQ=a (2) from 1 and 2 b+2PR=a b+2(b-c)=a b+2b-2c=a 3b-2c=a 3b=a+2c answer (D)